OpenStudy (anonymous):

The Golds' home was assessed this year at a value of $162,000. Last year it had been assessed at $150,000. What was the percent of increase?

OpenStudy (anonymous):

I think i know what u have to do

OpenStudy (anonymous):

try 150,000x = 162000 then divide by 150000 to isolate x. Then move decimal place two to the left. Not sure though

OpenStudy (texaschic101):

percent increase = (the new amount - the old amount)/ the old amount. Once you get that, multiply it by 100 to get the percent. I will set it up for you : percent increase = (162,000 - 150,000)/ 150,000 = x Now multiply : 100 * x = percent increase